OpenStudy (anonymous):

if dy/dx = 1+y, and when x=0, y=2, show that y = 3e^x -1

myininaya (myininaya):

use separation of variables cross multiply \[dy=(1+y)dx\] divide both sides by 1+y \[\frac{1}{1+y}dy=dx\] integrate both sides \[\int\limits_{}^{}\frac{1}{1+y}dy=\int\limits_{}^{}dx\] \[\ln|y+1|=x+C\] \[\ln|2+1|=0+C\] \[C=\ln(3)\] \[\ln|y+1|=x+\ln(3)\] lets solve this for y \[e^{\ln|y+1|}=e^{x+\ln3}\] \[y+1=e^x+e^{\ln3}\] \[y=e^x+3-1=e^x+2\]

myininaya (myininaya):

oops made an error e^(x+ln3) does not equal e^x+e^ln3 lol

myininaya (myininaya):

\[y+1=e^x*e^{\ln3}\] \[y+1=e^x*(3)\] \[y=3e^x-1\]

OpenStudy (anonymous):

thank you math genius!

OpenStudy (anonymous):

btw where did x+c come from?

myininaya (myininaya):

after you integrate you add constant

OpenStudy (anonymous):

integrate what?

myininaya (myininaya):

we integrated both sides the x came from integrating 1 with respect to x

myininaya (myininaya):

i got all my x's on one side and all my y's on the other side then i integrated to find solve both sides

OpenStudy (anonymous):

y+1=e^x*e^{ln3} ?

OpenStudy (anonymous):

i don't get that part as well

myininaya (myininaya):

law of exponents

myininaya (myininaya):

x^(a+b)=x^a*x^b

OpenStudy (anonymous):

y+1=e^x*(3) how did e^{ln3} become (3)? and what happened to the absolute value brackets around y+1